OpenStudy (anonymous):

the perimeter of a square is 60 units. The lengths of the sides of a square are 5x-2 and 3x+8..find x

OpenStudy (anonymous):

perimeter of a square = 4* side length

OpenStudy (ash2326):

If it's a square then the sides would be equal, can you compare the two sides?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I am just confused b/c my teacher did the problem like 5x-2+3x+8=60 but i thought that you need to do 5x-2+5x-2+3x+8+3x+8=60

OpenStudy (ash2326):

did the teacher do like this? \[2(5x-2+3x-8)=60\] I'm wondering, if this is a square then \[5x-2=3x-8\]

OpenStudy (anonymous):

it is a scaure and no he did not

OpenStudy (anonymous):

square

OpenStudy (ash2326):

Find x from the equation you wrote

OpenStudy (anonymous):

if it's a square, then \[5x-2=3x+8=60\]

OpenStudy (anonymous):

sorry, =15, not =60

OpenStudy (anonymous):

but i thought you double it??

OpenStudy (anonymous):

and y 15 ugg

OpenStudy (anonymous):

if the perimeter of a square is 60, the length of one of the sides is 15 because all the sides are the same length

OpenStudy (anonymous):

oo